OpenStudy (anonymous):

How do you convert 15 seconds into hrs.?

OpenStudy (ajprincess):

1h=3600s xh=15s x=15/3600

OpenStudy (anonymous):

One thing you might want to know when doing unit conversion is that you can use dimensional analysis, treating units like variables and multiply by 1. \[15 s = 15 s (1) = 15s(\frac{1hr}{60 min})(1)=15s(\frac{1hr}{60 min})(\frac{1min}{60s})=\frac{15}{3600}\]

OpenStudy (anonymous):

By the factor label method. There are 60 seconds in a minute and 60 minutes in an hour. 15 seconds x (1 minute/60 seconds) x (1 hour/60 minutes) = your answer Soon enough you'll realize that 60x60=3600 and so there are 3600 seconds in 1 hour. Then you will only need 1 conversion ratio; 15 seconds x (1 hour/3600 seconds) = your answer